# Semantics Prompting Data-Free Quantization for Low-Bit Vision Transformers

### v2

@@ Abstract
Data-free quantization (DFQ), which facilitates model quantization without real data to address increasing concerns about data security, has garnered significant attention within the model compression community. Recently, the unique architecture of vision transformers (ViTs) has driven the development of specialized DFQ techniques. However, we observe that the synthetic images from existing methods suffer from the deficient semantics issue compared to real images, thereby compromising performance. Motivated by this, we propose SPDFQ, a Semantics Prompting Data-Free Quantization method for ViTs. First, SPDFQ incorporates Attention Priors Alignment (APA), which uses randomly generated attention priors to enhance the semantics of synthetic images. Second, SPDFQ introduces Multi-Semantic Reinforcement (MSR), which utilizes localized patch optimization to prompt efficient parameterization and diverse semantics in synthetic images. Finally, SPDFQ employs Softlabel Learning (SL), where soft learning targets are adapted to encourage more complex semantics and accommodate images augmented by MSR. Experimental results demonstrate that SPDFQ significantly outperforms existing methods. For instance, SPDFQ achieves a 15.52% increase in top-1 accuracy on ImageNet for W4A4 ViT-B [^1] The code will be made publicly available upon acceptance..
Figure 1: Feature visualization using $t$-SNE [63]. Each shape (circle, star, triangle) represents a category. The red dashed circles highlight the features of our APA and real images. The visualized features are activations extracted before the classification head of DeiT-S. The features from PSAQ-ViT [43] deviate significantly from those of the real images, suggesting deficient semantics. While the features from our APA are more closely aligned with those of real images, suggesting improved semantics.
@@ 1 Introduction
Vision transformers (ViTs) [19] have attracted widespread interest from both academia and industry [33, 26] due to their superior performance across various vision tasks [62, 7, 1, 52]. However, the substantial computational demands and high memory requirements of ViTs pose significant challenges for deployment in resource-limited environments [61, 27, 29, 75, 42, 51, 11]. To address this challenge, model quantization [36] has emerged as a promising solution, reducing model complexity by enabling a low-bit representation of weights and activations. Nevertheless, most existing quantization methods rely on access to the original training data, raising concerns regarding data privacy and security. This data dependency restricts their application in data-sensitive environments [6, 77, 24, 70].
@@
To address this limitation, data-free quantization (DFQ) has been developed to quantize models without accessing real data, attracting considerable attention [69, 74, 14]. Most existing DFQ methods are tailored to convolutional neural networks (CNNs) and are not directly applicable to ViTs. These methods typically utilize batch normalization statistics (BNS), which capture the distribution of real data, to synthesize in-distribution synthetic data [6, 69, 74, 77]. However, BNS is unavailable for ViTs that use layer normalization (LN) to dynamically compute distribution statistics during inference [43]. Recently, several works have been proposed to accommodate the unique structure of ViTs [43, 44, 58, 16, 30]. For example, PSAQ-ViT [43] introduces patch similarity entropy (PSE) loss to optimize Gaussian noise towards usable synthetic images.
Nevertheless, we observe that the current synthetic images suffer from the deficient semantics issue. In Fig. 1, we visualize the penultimate features of the synthetic images using t-SNE [63]. It can be observed that the features of synthetic images generated by PSAQ-ViT diverge significantly from those of real images, suggesting that these synthetic images are semantically deficient. This observation is further supported by the quantitative results in Tab. 1, where the cosine similarity between synthetic images from PSAQ-ViT and real images is notably low. For high-bit quantization, where model capacity is largely retained [39], the performance degradation remains relatively minor even with deficient semantics images [43, 58]. However, in low-bit quantization, where model capacity is severely damaged and informative images are essential for recovering performance [77, 13], fine-tuning on these deficient semantics images leads to poor generalization to real datasets, resulting in limited performance gains.
@@
Motivated by this finding, we propose a Semantics Prompting Data-Free Quantization method for ViTs, termed SPDFQ. The overall framework is depicted in Fig. 2. First, SPDFQ introduces Attention Priors Alignment (APA), which employs randomly generated priors to encourage high responses in attention maps, thereby explicitly promoting semantics in synthetic images. These priors are generated using the Gaussian Mixture Model (GMM) to simulate diverse attention patterns. APA effectively prompts semantics in synthetic images, as validated by both visualization and quantitative analyses. As shown in Fig. 1, features of APA exhibit a closer alignment to those of real images, indicating high semantics. The quantitative results in Tab. 1 further confirm that the semantics of APA are more similar to the real images. Additionally, SPDFQ incorporates Multi-Semantic Reinforcement (MSR) and Softlabel Learning (SL). MSR promotes efficient parameterization and diverse semantics in synthetic images by leveraging localized patch-wise optimization, while SL prompts more complex semantics in synthetic images and accommodates images augmented by MSR through the use of soft learning targets.
@@
Experimental results across various ViT models demonstrate that SPDFQ presents substantial performance improvements. For example, SPDFQ achieves a 15.52% increase in top-1 accuracy on the ImageNet dataset for the W4A4 ViT-B model.
@@ 2.2 Network Quantization
Data-Driven Quantization. Model quantization reduces the complexity of neural networks by replacing full-precision weight and activation with the low-bit format. Data-driven quantization can be roughly divided into two categories: quantization-aware training (QAT) and post-training quantization (PTQ). QAT is compute-heavy since it re-trains the quantized model with the full training data to retain performance [40, 42, 38, 22, 48, 20, 67]. PTQ perform quantization with a tiny dataset and a reduced time overhead, harvesting widespread attention [3, 39]. The specific architecture of ViTs, such as LayerNorm and the self-attention module, urges distinct PTQ methods compared to CNNs [49, 18, 47, 42, 21, 79]. For example, Liu et al[51] develop a ranking loss to maintain the relative order of the self-attention activation. Unfortunately, both QAT and PTQ involve the original training data, causing concerns about data privacy and security issues in data-sensitive scenarios.
Data-Free Quantization. DFQ quantizes models without accessing real data [24, 14, 41, 2, 13, 57, 37, 32, 15]. Most previous DFQ methods focus on CNN, where the BNS can be adopted as the regularization term [6, 74]. However, BNS is infeasible for ViTs built on the LN. Recently, few efforts have been explored to accommodate ViTs [43, 44, 58, 16, 30]. PSAQ-ViT [43] introduces the first DFQ method for ViTs. They discover that Gaussian noise yields homogeneous patches, while the real image yields heterogeneous patches. Thus, patch similarity entropy (PSE) loss is proposed to optimize the Gaussian noise towards real-like images by making them showcase heterogeneous patches. Based on PSAQ-ViT, PSAQ-ViT V2 [44] further introduces a adversarial learning strategy [23]. [58] incorporates contrastive learning and proposes an iterative generation-quantization PTQ-based DFQ method. [30] proposes a sparse generation method to remove noisy and hallucination backgrounds in synthetic images.
Figure 2: Framework of the proposed SPDFQ. Attention Priors Alignment (APA) employs randomly generated attention priors to explicitly promote semantics in synthetic images. Multi-Semantic Reinforcement (MSR) prompts the different parts of synthetic images with various semantics. Softlabel Learning (SL) leverages soft learning to prompt more complex semantics in synthetic images.
@@ 3.1.1 Quantizers
[…]
@@ 3.1.2 Data Synthesis
DFQ methods parameterize synthetic images and optimize them toward real-like images with a pre-trained full-precision model $F$. Given a image $\tilde{\bm{I}}$ initialized from Gaussian noise, the one-hot loss [68] is introduced to prompt label-related semantics:
@@
To perform DFQ for ViTs, PSAQ-ViT [43] proposes patch similarity entropy (PSE) loss. It first compute patch similarity $\bm{\Gamma}_{l}[i,j]=\frac{u_{i}\cdot u_{j}}{||u_{i}||||u_{j}||}$, where $u_{i},u_{j}$ are feature vectors of MHSA outputs in $l$-th block and |||||$\cdot$||| | ⋅ | | denotes the $l_{2}$ norm. Then, it estimates the density function $\hat{f_{l}}(x)=\frac{1}{Mh}\sum_{m=1}^{M}K\left(\frac{x-x_{m}}{h}\right)$, where $K(\cdot)$ is a normal kernel, $h$ is the bandwidth, and $x_{m}$ is the kernel center derived from $\bm{\Gamma}_{l}$. Finally, the PSE loss is defined as:
[…]
@@ 3.2 Observations
Although existing DFQ methods of ViTs have made some progress, a significant performance gap remains between synthetic and real images. For example, as shown in Tab. 2, the W4A4 ViT-B fine-tuned on real images yields 68.16%, whereas PSAQ-ViT only achieves 36.32%. To dive into a deeper analysis, we visualize the features of the synthetic images before the classification head using t-SNE [63] in Fig. 1. Generally, the penultimate feature represents the semantics of the input [72, 4, 54, 71].
It is evident that features of PSAQ-ViT diverge significantly from those of real samples, indicating that the synthetic images fail to capture the real semantics distribution. We refer to this phenomenon as the deficient semantics issue. In Tab. 1, we quantitatively measure the semantics using the average cosine similarity (ranging from $-1$ to $1$). We also report the intra-class similarity within real images as an approximate upper bound for comparison. These results further highlight the deficient semantics of PSAQ-ViT. For example, for Class 1, the intra-class similarity within real images is 0.68, whereas PSAQ-ViT only presents 0.44, suggesting a low similarity between synthetic and real images. Similarly, for Class 3, while the real images exhibit a similarity of 0.41, PSAQ-ViT achieves only 0.31.
Both visualization and quantitative results reveal a pronounced deficient semantics. Consequently, quantized models fine-tuned on these semantically deficient synthetic images cannot well generalize to the real dataset, resulting in limited performance.
@@
Table 1: Average cosine similarity of three randomly selected classes. For real images, the similarity is measured within the class itself, while for PSAQ-ViT and APA, the similarity is measured between synthetic and real images of the same class. The results show that APA achieves higher similarity than PSE loss, indicating improved semantics.
3.3 Solutions
In the following, we introduce the proposed SPDFQ to prompt semantics in synthetic images. As illustrated in Fig. 2, SPDFQ introduces three contributions: Attention Priors Alignment (APA), Multi-Semantic Reinforcement (MSR), and Softlabel Learning (SL).
@@ 3.3.1 Attention Priors Alignment
The first step of SFDFQ in prompting semantics leverages the self-attention mechanism of ViTs. Specifically, the self-attention mechanism extracts and integrates semantic information by allocating more attention to tokens rich in semantics and less to tokens with minimal semantics [19, 28, 10]. Thus, if a synthetic image is semantically rich, its attention maps should exhibit high responses. Based on this characteristic, we propose Attention Priors Alignment (APA), which is designed to promote the semantics of synthetic images by explicitly encouraging high responses in images’ attention maps.
For a synthetic image $\tilde{\bm{I}}$, we first extract its attention maps of the $h$-th head in the $l$-th block, denoted by $\mathbf{A}_{l,h}\in\mathbb{R}^{N\times N}$, by inputting the image into a pre-trained full-precision model, where $N$ represents the total number of tokens. In DeiT, the attention of the classification token toward other tokens serves as the indicator for semantic versus non-semantic parts [5]. Accordingly, we extract $\mathbf{A}^{c}_{l,h}\in\mathbb{R}^{1\times(N-1)}$ from $\mathbf{A}_{l,h}$, representing the attention of the classification token to all tokens except itself. We then randomly generate attention priors $\tilde{\mathbf{A}}_{l,h}$, whose generation is detailed in the next part, and align $\mathbf{A}^{c}_{l,h}$ with $\tilde{\mathbf{A}}_{l,h}$ by:
@@
$\mathcal{L}^{\text{APA}}(\tilde{\bm{I}})=\sum_{l=S}^{L}\sum_{h=1}^{H}\frac{l}{L}\mathcal{L}_{l,h}(\tilde{\bm{I}}),$ (7)
@@
The proposed APA loss encourages synthetic images to contain semantically meaningful regions with high responses in attention maps, thereby prompting the semantics of the generated images.
@@
Attention Priors Generation. To generate attention priors $\tilde{\mathbf{A}}_{l,h}$, we employ GMM for it is the most commonly used distribution with high flexibility. ViTs utilize different attention heads to capture diverse patterns and learn varied, informative representations [8]. Thus, we use distinct GMMs for each head. Note that the goal here is not to replicate real attention maps precisely, but to generate simulated attention priors. These priors are used to explicitly encourage semantically meaningful regions within synthetic images, prompting high responses in attention maps to enhance semantics.
In particular, we first initialize an all zero matrix $\tilde{\mathbf{P}}\in\mathbb{R}^{H\times W}$, where $H=W=\sqrt{N-1}$ for DeiT, $H=W=\sqrt{N}$ for Swin. For example, for DeiT-S, the $H=W=\sqrt{196}=14$. Then, we generate $k$ two-dimensional Gaussian distributions, where $k$ is randomly sampled from $1\sim K_{APA}$ and $K_{APA}$ is set to $5$ in all experiments. Each Gaussian has its mean and covariance [^2] The pseudo code is detailed in supplementary materials.. Consequently, the matrix element at the $i$-th row and $j$-th column, $\tilde{\mathbf{P}}[i,j]$, is determined by:
@@
$\tilde{\mathbf{P}}_{n}=\frac{\tilde{\mathbf{P}}}{\sum\tilde{\mathbf{P}}}*(1-x).$ (9)
@@
Fig. 3 displays examples of generated attention priors. Although simple, APA effectively enhances semantics, as validated by both visualization and quantitative results. As shown in Fig. 1, compared to PSAQ-ViT, features after applying the APA loss are closer to the real images, indicating improved semantics. The quantitative results in Tab. 1 further support that APA achieves higher semantics. For instance, in Class 1, the intra-class similarity of PSAQ-ViT is 0.44, whereas APA achieves an improved similarity of 0.64.

### v5

# Semantic Alignment and Reinforcement for Data-Free Quantization of Vision Transformers

@@ Abstract
Data-free quantization (DFQ) enables model quantization without accessing real data, addressing concerns regarding data security and privacy. With the growing adoption of Vision Transformers (ViTs), DFQ for ViTs has garnered significant attention. However, existing DFQ methods exhibit two limitations: (1) semantic distortion, where the semantics of synthetic images deviate substantially from those of real images, and (2) semantic inadequacy, where synthetic images contain extensive regions with limited content and oversimplified textures, leading to suboptimal quantization performance. To address these limitations, we propose SARDFQ, a novel Semantics Alignment and Reinforcement Data-Free Quantization method for ViTs. To address semantic distortion, SARDFQ incorporates Attention Priors Alignment (APA), which optimizes synthetic images to follow randomly generated structure attention priors. To mitigate semantic inadequacy, SARDFQ introduces Multi-Semantic Reinforcement (MSR), leveraging localized patch optimization to enhance semantic richness across synthetic images. Furthermore, SARDFQ employs Soft-Label Learning (SL), wherein multiple semantic targets are adapted to facilitate the learning of multi-semantic images augmented by MSR. Extensive experiments demonstrate the effectiveness of SARDFQ, significantly surpassing existing methods. For example, SARDFQ improves top-1 accuracy on ImageNet by 15.52% for W4A4 ViT-B [^1] The code is at https://github.com/zysxmu/SARDFQ..
(a) The t-SNE[68] visualization of the penultimate-layer features (extracted by DeiT-S) of synthetic images. Each marker (circle, star, triangle) represents a distinct category. The red dashed circles highlight the features extracted from our APA and real images. Notably, the features produced by PSAQ-ViT[47] exhibit substantial deviation from those of real images, indicating semantic distortion. In contrast, our APA yields features that more closely align with those of real images, suggesting improved semantics alignment.
(b) The images of PSAQ-ViT and PSAQ-ViT V2 exhibit numerous dull regions with limited content and simplified textures, reflecting semantic inadequacy. In comparison, our SARDFQ generates images with greater diversity in both content and texture, demonstrating enhanced semantics.
Figure 1: Illustration of (a) semantic distortion and (b) semantic inadequacy.
@@ 1 Introduction
Vision Transformers (ViTs) [20, 34, 27] have demonstrated remarkable success across various computer vision tasks [67, 7, 1, 56]. However, their high computational cost and substantial memory footprint hinder deployment in resource-constrained environments [66, 28, 30, 81, 46, 55, 11]. To address these limitations, quantization [39] has emerged as a promising solution, which reduces model complexity by converting full-precision weight and activations into low-bit representations.
@@
Traditional quantization methods typically require access to the original training dataset, which raises data privacy and security concerns [6, 83, 25, 76]. As a result, data-free quantization (DFQ) has gained increasing attention, allowing quantization without the need for real data [75, 80, 14]. However, most existing DFQ methods are designed specifically for convolutional neural networks (CNNs) and are not directly applicable to vision transformers (ViTs). These methods generally rely on batch normalization statistics (BNS), which capture the distribution of real data, to synthesize in-distribution synthetic data [6, 75, 80, 83]. Yet, BNS is unavailable for ViTs, which use layer normalization (LN) to dynamically compute distribution statistics during inference [47]. Recently, several DFQ methods have been proposed for ViTs [47, 48, 62, 16, 31]. For example, PSAQ-ViT [47] introduces patch similarity entropy (PSE) loss to optimize Gaussian noise towards usable synthetic images.
Nevertheless, we observe that existing methods suffer from semantic distortion and inadequacy issues. As shown in Fig. 1(a) [^2] We do not plot the results of images from PSAQ-ViT V2 as they lack specific categories., features of synthetic images generated by PSAQ-ViT deviate significantly from those of real images. Tab. 1 shows that the cosine similarity between synthetic images from PSAQ-ViT and real images is notably low, also indicating significant distortion. These results highlight the issue of semantic distortion. Moreover, as shown in Fig. 1(b) [^3] More results are presented in the appendix., synthetic images generated by PSAQ-ViT and PSAQ-ViT V2 exhibit many regions with limited content diversity and overly simplified textures. These low-quality dull regions are useless or even detrimental to model learning [31], highlighting the issue of semantic inadequacy. Consequently, quantized models trained on such low-quality images suffer from degraded performance.
@@
Motivated by the above analysis, we propose a novel Semantics Alignment and Reinforcement Data-Free Quantization method for ViTs, termed SARDFQ. The overall framework is depicted in Fig. 2. To address the semantic distortion issue, SARDFQ introduces Attention Priors Alignment (APA), where synthetic images are optimized to follow structured attention priors generated using Gaussian Mixture Models (GMMs). APA effectively aligns the semantics of synthetic images with real images, as validated by both visual and quantitative analyses. As shown in Fig. 1(a), features of APA exhibit a closer alignment to those of real images, while quantitative results in Tab. 1 also confirm that the semantics of APA are more consistent to the real images, indicating enhanced semantics alignment. To address the semantic inadequacy issue, SARDFQ incorporates Multi-Semantic Reinforcement (MSR) and Softlabel Learning (SL). MSR utilizes localized patch optimization, which encourages different sub-patches of synthetic images to capture various semantics, reinforcing the rich semantics across images. SL applies multiple semantic targets to accommodate the learning of multi-semantic images augmented by MSR. As shown in Fig. 1(b), synthetic images after applying MSR exhibit greater diversity in content and texture, providing reinforced semantics.
@@
Experimental results across various ViT models and tasks demonstrate that SARDFQ presents substantial performance improvements. For example, SARDFQ achieves a 15.52% increase in top-1 accuracy on the ImageNet dataset for the W4A4 ViT-B model.
@@ 2.1 Vision Transformers
Figure 2: SARDFQ Framework overview: Attention Priors Alignment (APA) employs randomly generated attention priors to improve semantics alignment. Multi-Semantic Reinforcement (MSR) learns the different regions of synthetic images with various semantics to enhance overall semantic richness. Meanwhile, Softlabel Learning (SL) adopts multiple semantic targets to ensure consistent learning of multi-semantic images augmented by MSR.
@@ 2.2 Network Quantization
Data-Driven Quantization. Model quantization reduces the complexity of neural networks by replacing full-precision weight and activation with the low-bit format. Data-driven quantization can be roughly divided into two categories: quantization-aware training (QAT) and post-training quantization (PTQ). QAT is compute-heavy since it re-trains the quantized model with the full training data to retain performance [43, 46, 41, 23, 52, 21, 73, 65]. PTQ perform quantization with a tiny dataset and a reduced time overhead, harvesting widespread attention [3, 42]. The specific architecture of ViTs, such as LayerNorm and the self-attention module, urges distinct PTQ methods compared to CNNs [53, 18, 51, 46, 22, 85]. For example, Liu et al[55] develop a ranking loss to maintain the relative order of the self-attention activation. Unfortunately, both QAT and PTQ involve the original training data, causing concerns about data privacy and security issues in data-sensitive scenarios.
Data-Free Quantization. DFQ quantizes models without accessing real data [25, 14, 44, 2, 13, 61, 40, 33, 15, 45]. Most previous DFQ methods focus on CNN, where the BNS can be adopted as the regularization term [6, 80]. However, BNS is infeasible for ViTs built on the LN. Recently, few efforts have been explored to accommodate ViTs [47, 48, 62, 16, 31]. PSAQ-ViT [47] introduces the first DFQ method for ViTs. They discover that Gaussian noise yields homogeneous patches, while the real image yields heterogeneous patches. Thus, patch similarity entropy (PSE) loss is proposed to optimize the Gaussian noise towards real-like images by making them showcase heterogeneous patches. Based on PSAQ-ViT, PSAQ-ViT V2 [48] further introduces an adversarial learning strategy [24]. [62] incorporates contrastive learning and proposes an iterative generation-quantization PTQ-based DFQ method. [31] proposes a sparse generation method to remove noisy and hallucination backgrounds in synthetic images. SynQ [37] introduces an overall framework for DFQ by integrating a low-pass filter, class activation map alignment, and difficult samples learning. For a more comprehensive DFQ survey, we invite the reader to see [36].
@@ 3.1.1 Quantizers
where $\left\lfloor\cdot\right\rceil$ denotes rounding to the nearest integer, and clip limits the value to $0,2^{b}-1]$. Here, $\Delta$ and $z$ are the scale factor and zero-point, respectively. For the log2 quantizer:
@@ 3.1.2 Data Synthesis
DFQ methods parameterize synthetic images and optimize them toward real-like images with a pre-trained full-precision model $F$. Given a image $\tilde{\bm{I}}$ initialized from Gaussian noise, the one-hot loss [74] is introduced to learm label-related semantics:
@@
$\mathcal{L}^{\text{TV}}(\tilde{\bm{I}})=\iint|\nabla\tilde{\bm{I}}(\tau_{1},\tau_{2})|d\tau_{1}d\tau_{2}.$ (4)
@@
$\mathcal{L}^{\text{PSE}}(\tilde{\bm{I}})=\sum_{l=1}^{L}\int\hat{f_{l}}(x)\log\left[\hat{f_{l}}(x)\right]dx,$ (5)
@@ 3.2 Observations
Table 1: Average cosine similarity of three randomly selected classes. For real images, the similarity is measured within the class itself, while for PSAQ-ViT and APA, the similarity is measured between synthetic and real images of the same class. The results show that APA achieves higher similarity than PSE loss, indicating aligned semantics.
Existing DFQ methods have made significant progress. However, through carefully analyzing the synthetic images, we reveal that these images suffer from issues of semantic distortion and semantic inadequacy, both of which hinder further advancement in the DFQ of ViTs.
The semantic distortion issue refers to the significant divergence between the semantics of synthetic images and real images. To demonstrate this, we visualize the features of synthetic and real images in Fig. 1(a). Note that the penultimate feature typically is regarded to represent the semantics of the input [78, 4, 58, 77], It is clear that the features of PSAQ-ViT diverge significantly from real images, suggesting that these images fail to capture the true semantic distribution of real data. Tab. 1 quantitatively measures the semantics using the average cosine similarity (ranging from 1-1 to $1$). We also report the intra-class similarity within real images as an approximate upper bound for comparison. This result further supports the observation of low similarity between synthetic images from PSAQ-ViT and real images. For instance, for Class 1, the intra-class similarity within real images is 0.68, while PSAQ-ViT achieves only 0.44.
The semantic inadequacy issue refers to the presence of dull regions in synthetic images, which contain redundant or non-semantic content [31, 35], hindering the model’s learning process. As indicated in [14, 13], a diverse content and textures generally suggests rich information. As shown in Fig. 1(b), many regions of synthetic images generated by PSAQ-ViT and PSAQ-ViT V2 exhibit a lack of diversity in content, with overly simplified textures. Specifically, the central region of PSAQ-ViT images only contains faint object structures, while PSAQ-ViT V2 images appear excessively smoothed and indistinct.
For high-bit quantization, where model capacity is largely retained [42], the performance degradation remains relatively minor even using semantic distorted and inadequate images [47, 62]. However, in low-bit quantization, where model capacity is severely damaged and informative images are essential for recovering performance [83, 13], fine-tuning on these poor-quality images leads to inferior generalization to real datasets, resulting in limited performance. For example, as shown in Tab. 2, the W4A4 ViT-B fine-tuned on real images yields 68.16%, whereas PSAQ-ViT only achieves 36.32%.
Figure 3: Comparison between attention maps.
3.3 Semantics Alignment and Reinforcement Data-Free Quantization
In the following, we introduce the proposed SARDFQ, whose framework is illustrated in Fig. 2.
@@ 3.3.1 Attention Priors Alignment
In ViTs, the self-attention mechanism encodes semantic correlations between image regions, where high-response areas in attention maps strongly correlate with semantic-discriminative content [20, 29, 10]. However, existing DFQ methods overlook this intrinsic property in the generation process. As a result, as shown in Fig. 3, synthetic images often exhibit disordered and unnatural attention patterns, with attention maps either overly diffuse or misaligned toward peripheral regions. This undermines their ability to preserve semantic-discriminative content, causing semantic distortion, as demonstrated in Fig. 1(a) and Tab. 1. In response, we propose Attention Priors Alignment (APA), which improves semantics alignment by optimizing synthetic images to follow randomly generated structure attention priors.
Specifically, given a synthetic image $\tilde{\bm{I}}$, we first obtain its attention maps in the $h$-th head of the $l$-th block, denoting as $\mathbf{A}_{l,h}\in\mathbb{R}^{N\times N}$, where $N$ represents the total number of tokens. In DeiT, the attention of the classification token toward other tokens serves as the indicator for semantic versus non-semantic parts [5]. Thus, we extract $\mathbf{A}^{c}_{l,h}\in\mathbb{R}^{1\times(N-1)}$ from $\mathbf{A}_{l,h}$, representing the attention of the classification token to all tokens except itself. We then randomly generate attention priors $\tilde{\mathbf{A}}_{l,h}$, whose generation is detailed in the next part, and align $\mathbf{A}^{c}_{l,h}$ with $\tilde{\mathbf{A}}_{l,h}$ by:
@@
[…]
@@
Attention Priors Generation. To generate attention priors $\tilde{\mathbf{A}}_{l,h}$, Gaussian Mixture Models (GMMs) are employed as it is the most commonly used distribution with high flexibility. ViTs utilize different attention heads to capture diverse patterns and learn varied, informative representations [8]. Thus, we use distinct GMMs for each head. Note that the goal here is not to replicate real attention maps precisely, but to generate simulated structure attention priors to guide the learning of synthetic images.
In particular, we first initialize an all zero matrix $\tilde{\mathbf{P}}\in\mathbb{R}^{H\times W}$, where $H=W=\sqrt{N-1}$ for DeiT, $H=W=\sqrt{N}$ for Swin. For example, for DeiT-S, the $H=W=\sqrt{196}=14$. Then, we generate $k$ two-dimensional Gaussian distributions, where $k$ is randomly sampled from $1\sim K_{APA}$ and $K_{APA}$ is set to $5$ in all experiments. Each Gaussian has its mean and covariance [^4] The pseudo code is detailed in appendix.. Consequently, the matrix element at the $i$-th row and $j$-th column, $\tilde{\mathbf{P}}[i,j]$, is determined by:
@@
$\tilde{\mathbf{P}}_{n}=\frac{\tilde{\mathbf{P}}}{\sum\tilde{\mathbf{P}}}\cdot(1-x).$ (9)
@@
Fig. 4 displays examples of generated attention priors.
Discussion. APA prevents attention disorder, ensuring that synthetic images exhibit more coherent and natural attention patterns, as demonstrated in Fig. 3. As a result, APA selectively enhances the responses in certain regions, effectively emphasizing semantic-discriminative regions within synthetic images and thus prompting the discriminative features. Although simple, APA enables synthetic images to align better with the real semantics, as validated by both visual and quantitative evaluations. As shown in Fig. 1(a), compared to PSAQ-ViT, features obtained after applying the APA loss are more closely aligned with real images, indicating better semantic alignment. The quantitative results in Tab. 1 further support that APA achieves superior semantic alignment. For example, in Class 1, the intra-class similarity for PSAQ-ViT is 0.44, whereas APA achieves a higher similarity of 0.64.